OpenStudy (anonymous):

In a circle of radius 10 cm, a sector has an area of 40(pi) sq. cm. What is the degree measure of the arc of the sector?

OpenStudy (anonymous):

pi r ^2 = 100 pi

OpenStudy (anonymous):

with r = 10

OpenStudy (anonymous):

so it's 40/100

OpenStudy (anonymous):

40/100 = x/360

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

360 * 40

OpenStudy (anonymous):

then divide by 100 then you get your angle

OpenStudy (anonymous):

144?

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

1.44?

OpenStudy (anonymous):

you just take what you know.. put it in a formula and solve

OpenStudy (anonymous):

no 144

OpenStudy (anonymous):

360 times 40 = 14400

OpenStudy (anonymous):

then divide by 100

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

Thank you again!